# PlaceFM: A Training-free Geospatial Foundation Model of Places using Large-Scale Point of Interest Data

### v2

@@ 4.1. FSQ-19M POI Dataset Preparation
Pre-processing pipeline: After extracting all POIs located within the United States, we performed several pre-processing and enhancement steps to construct FSQ-19M. For each POI $p$, we retained its latitude ($\text{lat}_{p}$), longitude ($\text{lon}_{p}$), state, locality (city), and ZIP code. These attributes provide the spatial anchor for each point.
Table 1. A data entry in the FSQ-19M dataset.
Attribute Value
longitude 40.858911
latitude -124.073245
state CA
locality Arcata
postcode 95521
date_created 2012-02-08
date_closed None
Each POI is associated with the date it was created and, if applicable, the date it was closed, enabling temporal analysis of the evolution of urban spaces. Also, each POI is assigned a hierarchical semantic category, defined up to six levels. This hierarchical structure allows for flexible aggregation at different levels of semantic granularity. We applied preprocessing steps to normalize textual attributes (e.g., consistent naming of states and ZIP codes) and to remove duplicates. Moreover, we performed a spatial join with U.S. Census ZIP code boundary data to validate and, if necessary, correct the location of each POI. The enriched dataset covers diverse urban functions, dominated by dining venues such as pizzerias, coffee shops, and fast food restaurants. Fuel stations, hair salons, churches, and fire stations are also well represented, reflecting the dataset’s commercial and civic relevance for various spatial and socioeconomic analyses.
After preprocessing, the dataset consists of 19,019,187 POIs distributed across all 48 contiguous states in the U.S., each with spatial, temporal, and semantic information. Figure 1 visualizes the spatial distribution of POIs using a uniform random sample of one million entries from FSQ-19M. Also, Table 1 illustrates a representative sample from FSQ-19M, showing both spatial and semantic attributes. Section 4.2 describes the methodology used to construct a graph structure from the FSQ-19M POI dataset.
Table 2. General statistics of the FSQ-19M POI graph dataset
19,019,187 44,202,952 48 31,970 60,155
Figure 3. Graph structures of a sample of POIs within ZIP code 10031 in Manhattan, NY: (a) 4-NN graph (unweighted), (b) 8-NN graph (unweighted), and (c) region-adaptive Delaunay triangulation (weighted). Edge opacity reflects weight, with darker edges indicating stronger connections.
4.2. POI Graph Data Construction
@@
By representing POIs as nodes in a graph, each place, defined as a POI together with its surrounding spatial context, can be effectively modeled through the message-passing mechanism in GCNs. This allows each POI embedding to be enriched by propagating information from its neighboring POIs, thereby capturing the contextual semantics that shape the uniqueness of a place. The advantage of adopting a graph representation is that it provides a flexible and compact structure for modeling POIs, remains robust to spatial transformations, and naturally encodes contextual dependencies through message passing (Huang et al., 2023). Various strategies exist to construct such POI graphs, depending on the choice of proximity functions or similarity metrics for edge formation. In this study, we adopt two widely used strategies for constructing the graph structure based on geographic proximity.
@@
Here, $L_{r}$ is a region-specific scaling factor, computed using the POIs within the subgraph $G_{r}\subseteq G_{s}$ corresponding to a local region of POIs. This local, adaptive scaling is important because the spatial density of POIs can vary drastically even within the same state: for instance, POIs in a dense urban subarea such as Manhattan, NY, are much closer together than POIs in a less dense area of upstate New York, such as the Adirondack region. By computing $L_{r}$ for each region $G_{r}$, the weighting function remains balanced, ensuring that the edge weights are comparable across areas with highly variable local densities. Following (Huang et al., 2023), the factor $w_{r}(i,j)$ encodes regional consistency: it is set to $1.0$ if POIs $i$ and $j$ belong to the same region, and to $0.4$ if they connect across different regions, thereby down-weighting cross-region connections. The computed edge weights are then scaled uniformly within each region to the range $[0,1]$, resulting in $\mathbf{A}_{s}$ being a normalized weighted adjacency matrix.
(2) k-Nearest Neighbors (KNN). For each POI $i$, we connect its $k$ closest neighbors based on geographic distance (Euclidean or Haversine) in the adjacency matrix $\mathbf{A}_{s}$. Edges are unweighted (0 or 1), indicating the presence of a connection. Despite its simplicity, this approach balances neighbor consistency and spatial proximity: each node connects to its nearest neighbors, capturing local interactions while preventing overly dense graphs in high-density areas. Its efficiency and binary encoding make KNN graphs a practical choice for large-scale POI datasets and downstream GNN operations.
Figure 3 illustrates an example of different POI connectivity patterns in the 10031 ZIP code of Manhattan, NY. After constructing the graph structures for FSQ-19M, the statistics for all 48 graphs across the contiguous United States are summarized in Table 2.
4.3. Training-free Place Foundation Model
PlaceFM is a training-free foundation model of places, designed to generate meaningful urban region embeddings from graph-based POI data and to identify places within a given geographical region. The overall architecture of PlaceFM is shown in Figure 2. In this section, we provide a detailed description of each component of the framework:
4.3.1. Feature Encoder
Our proposed PlaceFM framework enables encoding POI-level raw attributes to extract rich and meaningful features, facilitating effective representation learning for a variety of urban-related tasks. In this study, we leverage the semantic category of each POI as the primary attribute, which is inherently hierarchical, capturing fine-grained distinctions between different types of POIs. To model these hierarchical category attributes, we utilize SD-CEM (Jia et al., 2024), a semantically disentangled POI category embedding model. It generates hierarchy-enhanced category representations by pre-training on large-scale mobility sequences, learning disentangled embeddings that capture semantic relationships among POI categories.
@@ 4.3.2. POI Encoder
To capture the contextual semantics of places, it is therefore essential to move beyond isolated POI attributes and incorporate region-level structural information. For instance, a Starbucks located near university buildings and departments carries very different contextual meaning compared to one situated along a remote highway rest stop. To this end, PlaceFM employs a lightweight, non-parametric graph propagation mechanism that enriches each POI’s representation by incorporating information from its spatial neighbors.
@@
(2) $\mathbf{H}^{(k)}=\hat{\mathbf{A}}^{k}\mathbf{V},\quad\mathbf{V}\in\mathbb{R}^{|\mathcal{V}|\times d},$
where $\hat{\mathbf{A}}$ is the symmetrically normalized adjacency matrix of $G$, defined as:
(3) $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}},$
@@
(4) $\mathbf{H}=\sum_{k=0}^{K}\alpha_{k}\mathbf{H}^{(k)},$
where $\alpha_{k}\in\mathbb{R}$ is the propagation weight at step $k$. This formulation controls the extent to which semantic and spatial information is propagated, thereby allowing PlaceFM to act as a spatially and semantically controlled approach. In doing so, it ensures that each POI’s embedding integrates both local and multi-hop neighborhood semantics. Importantly, we allow $\alpha_{k}$ to vary across a wide range, including negative values. When $\alpha_{k}<0$, the model effectively introduces a “negative offset” that captures heterophilic relationships between POIs (Zhu et al., 2021), which is particularly useful in urban settings where co-located POIs may have contrasting functions (e.g., a bar next to a church).
4.3.3. Place Representation Learning via Clustering
Once we obtain the POI embeddings, the next step is to generate region-level embeddings and to identify places within a region, as introduced as the two main objectives of PlaceFM in Section 3. Unlike existing geospatial foundation models for urban representation learning, which are typically restricted to a fixed spatial resolution, PlaceFM is inherently multigranular, enabling the generation of embeddings at arbitrary levels of granularity. This flexibility accommodates both geographic zones (e.g., ZIP codes, cities, census block groups) and non-geographic groupings (e.g., functional regions, semantic clusters).
From a data-centric AI perspective (Zha et al., 2025), one effective approach to summarizing a large-scale POI graph while preserving its semantic and spatial structure is to learn representative embeddings for aggregated POIs within each region. This process, known as graph reduction (Hashemi et al., 2024), reduces the size of the graph while retaining sufficient expressivity for downstream tasks. Inspired by GECC (Gong et al., 2025), a clustering-based graph reduction method that guarantees comparable performance for GNNs trained on condensed graphs, PlaceFM introduces an urban representation learning component that leverages clustering over propagated POI embeddings to construct semantically meaningful and context-aware places. This transforms fine-grained POI embeddings into higher-level place representations.
@@ 4.3.3. Place Representation Learning via Clustering
(5) $\mathbf{H}_{r}=\{\mathbf{h}_{p}\mid p\in\mathcal{V}_{r}\},\quad\mathbf{H}_{r}\in\mathbb{R}^{n_{r}\times d},$
@@
(6) $\min_{\{\mathcal{C}_{1},\dots,\mathcal{C}_{k_{r}}\}}\sum_{j=1}^{k_{r}}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\|\mathbf{h}_{p}-\boldsymbol{\mu}_{j}\|_{2}^{2},$
@@
Following GECC, we introduce a reduction ratio hyperparameter $r\in[0,1]$ to control the number of clusters per region: $k_{r}=\left\lfloor n_{r}\times r\right\rfloor$, where smaller values of $r$ yield fewer but coarser places, and larger values of $r$ produce finer-grained places. This formulation allows flexible trade-offs between representational detail and computational efficiency. In cases where the distribution of POIs is highly non-uniform, density-based clustering methods such as DBSCAN (Ester et al., 1996) can also be employed, which do not require specifying $k_{r}$ in advance.
@@
The resulting set of clusters $\{\mathcal{C}_{1},\dots,\mathcal{C}_{k_{r}}\}$ serves as the places representing region $r$. Each cluster is summarized by its centroid embedding:
(7) $\mathbf{z}_{j}=\frac{1}{|\mathcal{C}_{j}|}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\mathbf{h}_{p},\quad j=1,\dots,k_{r},$
@@
yielding a condensed representation $\mathbf{Z}_{r}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{k_{r}}\}\in\mathbb{R}^{k_{r}\times d}$. It has been theoretically proven (Gong et al., 2025) that such condensed representations are as expressive as the original POI embeddings, allowing downstream models trained on $\mathbf{Z}_{r}$ to achieve comparable performance to those trained on $\mathbf{H}_{r}$, while the size of $\mathbf{Z}_{r}$ is significantly smaller than $\mathbf{H}_{r}$, resulting in much higher efficiency during model training (Hashemi et al., 2024).
4.3.4. Region Aggregator
@@
Once the place embeddings $\mathbf{Z}_{r}$ are obtained, in order to generate a single meaningful representation for the entire region $r$, we perform a simple and efficient aggregation function agg, which computes the weighted average of all place embeddings, weighted by the number of POIs in each place. Formally, the aggregated regional embedding is given by: $\mathbf{L}_{r}=\big[\,\mathbf{l}_{1}\;|\;\mathbf{l}_{2}\;|\;\dots\;|\;\mathbf{l}_{N_{r}}\,\big]$, where each $\mathbf{l}_{j}$ is a place embedding computed as the weighted average of POIs in place $\mathbf{l}_{j}=\frac{\sum_{i=1}^{N_{j}}n_{i}\cdot\mathbf{z}_{i}}{\sum_{i=1}^{N_{j}}n_{i}}$, where $\mathbf{z}_{i}$ is the embedding of place $i$, $n_{i}$ is the number of POIs contained in place $i$, and $N_{r}$ is the total number of places in region $r$. This weighted aggregation works because it ensures that places with more POIs, reflecting denser activity or higher functional importance, contribute proportionally more to the regional representation. In other words, the embedding $\mathbf{l}_{r}$ captures both the semantic characteristics of places and the relative concentration of POIs, leading to a more faithful and efficient summary of the entire region. If needed, we can also construct an adjacency matrix to represent the connectivity between region embeddings, where two regions are considered connected if their polygon boundaries touch each other.
5. Experiments
@@
Table 3. ZIP-code level prediction performance for population density $\left(\text{people}/\text{km}^{2}\right)$. The best results are highlighted in bold, and values in parentheses indicate the standard deviation.
@@
To evaluate the effectiveness of PlaceFM, we compare it against the state-of-the-art baselines in urban representation learning. We begin by describing the experimental setup and implementation details of each method. We then report the performance of two urban downstream tasks: (i.) Population Density Prediction and (ii.) Housing Price Prediction, as measures of representation quality, along with a comparison of computational efficiency. Furthermore, we demonstrate the effectiveness of the identified places both quantitatively and qualitatively through case studies on selected regions using limited ground-truth data. Finally, we conduct an ablation study to empirically assess the contribution of each module in our framework.
5.1. Experimental Setup
5.1.1. Dataset and Baselines
@@
The FSQ-19M POI graph dataset is constructed as a collection of 48 homogeneous graphs, each corresponding to one of the contiguous U.S. states. Splitting the dataset at the state level allows us to evaluate PlaceFM across graphs with diverse densities and structural statistics. For instance, Wyoming (WY) forms the smallest graph, containing 40,165 POIs, while California (CA) constitutes the largest, with 2,204,300 POIs. This setup provides a natural testbed for assessing the scalability and robustness of PlaceFM in comparison with existing methods under varying graph sizes and complexities.
@@
We benchmark PlaceFM against several representative baselines, ranging from simple heuristics to state-of-the-art geospatial foundation models, all of which generate region-level embeddings within each state. For a fair comparison, the dimensionality of the generated region embeddings is fixed at $d=30$ across all models.
2. Place2Vec (Yan et al., 2017): A method that learns POI category embeddings and then derives region embeddings by averaging over the POIs contained in a region. Unlike simple averaging, Place2Vec incorporates spatial co-occurrence statistics during POI embedding learning, yet the final regional representation is still obtained through aggregation by averaging.
3. HGI (Huang et al., 2023): An unsupervised model that learns region embeddings by jointly modeling categorical semantics of POIs, POI-level and region-level adjacency, and the interaction between POIs and regions. HGI further employs an attention mechanism to weigh the relative importance of individual POIs during region-level aggregation, enabling more context-sensitive representations.
4. PDFM (Agarwal et al., 2024): A pre-trained foundation model that integrates diverse geospatial signals, including POI data, maps, activity levels, search trends, weather, and air quality, into a heterogeneous graph. Geographic regions such as counties and postal codes are represented as nodes, with edges reflecting spatial proximity. A GNN is then applied to capture spatial dependencies and generate embeddings for these regions.
5.1.2. Implementation Details
To ensure a fair reproduction and comparison with baseline methods, we reimplemented all baselines and tuned their hyperparameters, guided both by the best configurations reported in the original papers and by additional fine-tuning under our experimental setting. The complete implementation, including code for PlaceFM and all baseline models, is publicly available in the PlaceFM repository [^4] https://github.com/mohammadhashemii/PlaceFM. To maintain consistency, the number of evaluations for downstream tasks was restricted to 10. All baselines, with the exception of PDFM (Agarwal et al., 2024), were trained from scratch. For PDFM, due to the unavailability of detailed architectural specifications, we reused the released pretrained embeddings. Since PDFM integrates heterogeneous data sources (e.g., maps, activity levels, search trends, weather, and air quality), we only utilized the map-based embeddings and further applied dimensionality reduction to match the embedding size to $d=30$, ensuring a relatively fair comparison with other methods.
For PlaceFM, we determined that the optimal dimensionality of the feature encoder was $d=30$. The POI encoder was implemented using a single-layer SGC (Wu et al., 2019) with propagation limited to two steps (capturing up to second-order neighborhood information). The propagation coefficients were tuned within the range $\alpha_{0},\alpha_{1},\alpha_{2}\in[0.0,1.0]$ with increments of $0.25$. For the clustering module, the maximum number of iterations for $k$-means was set to 300, with a convergence threshold of $1\times 10^{-8}$. Each experiment was repeated 10 times, and we report the averaged results to ensure statistical reliability.
To efficiently execute the clustering procedure, we relied on Intel(R) Xeon(R) Platinum 8260 CPUs @ 2.40GHz with NumPy (Harris et al., 2020) for numerical computation. Baseline model training was conducted on a high-performance computing cluster equipped with a heterogeneous mix of GPUs: Tesla A100 (40GB) and V100 (32GB) for large-scale graphs, and K80 (12GB) GPUs for smaller graphs. This heterogeneous setup enabled efficient handling of datasets of varying sizes while ensuring consistent evaluation across all methods.
5.2. Downstream tasks performance analysis
5.2.1. Population Density Prediction

### v3

@@ 4.1 FSQ-19M POI Dataset Preparation
After extracting all POIs located within the United States, we performed several pre-processing and enhancement steps to construct FSQ-19M. The complete preprocessing and data enrichment pipeline used to construct FSQ-19M is described in detail in Appendix A.1. After preprocessing, the dataset consists of 19,019,187 POIs distributed across all 48 contiguous states in the U.S., each with spatial, temporal, and semantic information. See Appendix A for dataset details, statistics, and visualizations.
4.2 POI Graph Data Construction
@@
By modeling POIs as nodes in a graph, each place, defined as a POI and its surrounding spatial context, can be represented through message passing in GCNs, enriching POI embeddings with information from neighboring POIs and capturing contextual semantics. Graph representations provide a flexible and compact structure for modeling POIs, are robust to spatial transformations, and naturally encode contextual dependencies through message passing (Huang et al., 2023). In this study, we adopt two widely used strategies for constructing the graph structure based on geographic proximity.
@@
Here, $L_{r}$ is a region-specific scaling factor, computed using the POIs within the subgraph $G_{r}\subseteq G_{s}$ corresponding to a local region of POIs. This local, adaptive scaling is important because the spatial density of POIs can vary drastically even within the same state. By computing $L_{r}$ for each region $G_{r}$, the weighting function remains balanced, ensuring that the edge weights are comparable across areas with highly variable local densities. Following (Huang et al., 2023), the factor $w_{r}(i,j)$ encodes regional consistency: it is set to $1.0$ if POIs $i$ and $j$ belong to the same region, and to $0.4$ if they connect across different regions, thereby down-weighting cross-region connections. The computed edge weights are then scaled uniformly within each region to the range $[0,1]$, resulting in $\mathbf{A}_{s}$ being a normalized weighted adjacency matrix.
(2) k-Nearest Neighbors (KNN). For each POI $i$, we connect its $k$ closest neighbors based on geographic distance (Euclidean or Haversine) in the adjacency matrix $\mathbf{A}_{s}$. Edges are unweighted (0 or 1), indicating the presence of a connection. Despite its simplicity, this approach balances neighbor consistency and spatial proximity: each node connects to its nearest neighbors, capturing local interactions while preventing overly dense graphs in high-density areas.
4.3 Training-free Place Representation Generation
PlaceRep is a training-free method designed to generate meaningful urban region embeddings from graph-based POI data and to identify places within a given geographical region. The overall architecture of PlaceRep is shown in Figure 1. In this section, we provide a detailed description of each component of the framework:
4.3.1 Feature Encoder
PlaceRep enables encoding POI-level raw attributes to extract rich and meaningful features, facilitating effective representation learning for a variety of urban-related tasks. In this study, we leverage the semantic category of each POI as the primary attribute, which is inherently hierarchical, capturing fine-grained distinctions between different types of POIs. To model these hierarchical category attributes, we utilize SD-CEM (Jia et al., 2024), a semantically disentangled POI category embedding model. It generates hierarchy-enhanced category representations by pre-training on large-scale mobility sequences, learning disentangled embeddings that capture semantic relationships among POI categories.
@@ 4.3.2 POI Encoder
To capture the contextual semantics of places, it is therefore essential to move beyond isolated POI attributes and incorporate region-level structural information. For instance, a Starbucks located near university buildings and departments carries very different contextual meaning compared to one situated along a remote highway rest stop. To this end, PlaceRep employs a lightweight, non-parametric graph propagation mechanism that enriches each POI’s representation by incorporating information from its spatial neighbors.
@@
$\mathbf{H}^{(k)}=\hat{\mathbf{A}}^{k}\mathbf{V},\quad\mathbf{V}\in\mathbb{R}^{|\mathcal{V}|\times d},$ (2)
where $\hat{\mathbf{A}}$ is the symmetrically normalized adjacency matrix of $G$, defined as $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$.
@@
$\mathbf{H}=\sum_{k=0}^{K}\alpha_{k}\mathbf{H}^{(k)},$ (3)
where $\alpha_{k}\in\mathbb{R}$ is the propagation weight at step $k$. This formulation controls the extent to which semantic and spatial information is propagated, thereby allowing PlaceRep to act as a spatially and semantically controlled approach. In doing so, it ensures that each POI’s embedding integrates both local and multi-hop neighborhood semantics. Importantly, we allow $\alpha_{k}$ to vary across a wide range, including negative values. When $\alpha_{k}<0$, the model effectively introduces a “negative offset” that captures heterophilic relationships between POIs (Zhu et al., 2021), which is particularly useful in urban settings where co-located POIs may have contrasting functions (e.g., a bar next to a church).
4.3.3 Place Representation Learning via Clustering
Once we obtain the POI embeddings, the next step is to generate region-level embeddings and to identify places within a region, as introduced as the two main objectives of PlaceRep in Section 3. Unlike existing urban region representation learning methods, which are typically restricted to a fixed spatial resolution, PlaceRep is inherently multigranular, enabling the generation of embeddings at arbitrary levels of granularity. This flexibility accommodates both geographic zones and non-geographic groupings (e.g., functional regions, semantic clusters).
From a data-centric AI perspective (Zha et al., 2025), one effective approach to summarizing a large-scale POI graph while preserving its semantic and spatial structure is to learn representative embeddings for aggregated POIs within each region. This process, known as graph reduction (Hashemi et al., 2024), reduces the size of the graph while retaining sufficient expressivity for downstream tasks. Inspired by GECC (Gong et al., 2025), a clustering-based graph reduction method that theoretically guarantees comparable performance for GNNs trained on condensed graphs, PlaceRep introduces an urban representation learning component that leverages clustering over propagated POI embeddings to construct semantically meaningful and context-aware places. This transforms fine-grained POI embeddings into higher-level place representations.
@@ 4.3.3 Place Representation Learning via Clustering
$\mathbf{H}_{r}=\{\mathbf{h}_{p}\mid p\in\mathcal{V}_{r}\},\quad\mathbf{H}_{r}\in\mathbb{R}^{n_{r}\times d},$ (4)
@@
$\min_{\{\mathcal{C}_{1},\dots,\mathcal{C}_{k_{r}}\}}\sum_{j=1}^{k_{r}}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\|\mathbf{h}_{p}-\boldsymbol{\mu}_{j}\|_{2}^{2},$ (5)
@@
We introduce a reduction ratio hyperparameter $r\in[0,1]$ to control the number of places per region: $k_{r}=\left\lfloor n_{r}\times r\right\rfloor$, where smaller values of $r$ yield fewer but coarser places, and larger values of $r$ produce finer-grained places. This formulation allows flexible trade-offs between representational detail and computational efficiency. In cases where the distribution of POIs is highly non-uniform, density-based clustering methods such as DBSCAN (Ester et al., 1996) can also be employed, which do not require specifying $k_{r}$ in advance.
@@
The resulting set of clusters $\{\mathcal{C}_{1},\dots,\mathcal{C}_{k_{r}}\}$ serves as the places representing region $r$. Each cluster is summarized by its centroid embedding: $\mathbf{z}_{j}=\frac{1}{|\mathcal{C}_{j}|}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\mathbf{h}_{p}$, where
$\mathbf{z}_{j}=\frac{1}{|\mathcal{C}_{j}|}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\mathbf{h}_{p},\quad j=1,\dots,k_{r},$ (6)
@@
yielding a condensed representation $\mathbf{Z}_{r}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{k_{r}}\}\in\mathbb{R}^{k_{r}\times d}$. It has been theoretically proven (Gong et al., 2025) that such condensed representations are as expressive as the original POI embeddings, allowing downstream models trained on $\mathbf{Z}_{r}$ to achieve comparable performance to those trained on $\mathbf{H}_{r}$, while the size of $\mathbf{Z}_{r}$ is significantly smaller than $\mathbf{H}_{r}$, resulting in much higher efficiency during model training.
4.3.4 Region Aggregator
@@
Once the place embeddings $\mathbf{Z}_{r}$ are obtained, in order to generate a single meaningful representation for the entire region $r$, a simple and efficient aggregation function agg is performed, which computes the weighted average of all place embeddings, weighted by the number of POIs in each place. Formally, the aggregated regional embedding is given by: $\mathbf{L}_{r}=\big[\,\mathbf{l}_{1}\;|\;\mathbf{l}_{2}\;|\;\dots\;|\;\mathbf{l}_{N_{r}}\,\big]$, where each $\mathbf{l}_{j}$ is a place embedding computed as the weighted average of POIs in place $\mathbf{l}_{j}=\frac{\sum_{i=1}^{N_{j}}n_{i}\cdot\mathbf{z}_{i}}{\sum_{i=1}^{N_{j}}n_{i}}$, where $\mathbf{z}_{i}$ is the embedding of place $i$, $n_{i}$ is the number of POIs contained in place $i$, and $N_{r}$ is the total number of places in region $r$. An adjacency matrix is constructed to represent the connectivity between region embeddings, where two regions are considered connected if their polygon boundaries touch each other.
@@
Table 1: ZIP-code level prediction performance for population density $\left(\text{people}/\text{km}^{2}\right)$. The best results are highlighted in bold, and values in parentheses indicate the standard deviation.
State (# POIs) Averaging Place2Vec (Yan et al., 2017) HGI (Huang et al., 2023) PDFM (Agarwal et al., 2024) PlaceRep (ours)
@@
5 Experiments
5.1 Experimental Setup
@@
The FSQ-19M POI graph dataset is constructed as a collection of 48 homogeneous graphs, each corresponding to one of the contiguous U.S. states. Splitting the dataset at the state level allows us to evaluate PlaceRep across graphs with diverse densities and structural statistics. For instance, Wyoming (WY) forms the smallest graph, containing 40,165 POIs, while California (CA) constitutes the largest, with 2,204,300 POIs.
@@
We benchmark PlaceRep against several representative baselines, ranging from simple heuristics to state-of-the-art geospatial foundation and region representation learning models, all of which generate region-level embeddings within each state. See Appendix C.2 for details of these baselines.
5.2 Downstream tasks performance analysis
5.2.1 Population Density Prediction